An electromagnetic wave with a peak magnetic field of magnitude 1.50 x 10^(-7) T has an associated peak electric field of what magnitude?

A. 0.500 x 10^(-15) N/C

B. 2.00 x 10^(-5) N/C

C. 2.20 x 10^(4) N/C

D. 45.0 N/C

E. 22.0 N/C


1
Expert's answer
2021-02-18T18:47:32-0500

"Magnitude\:of\:peak\:electric\:field\left(E\right) =Velocity\:of\:light\:\left(C\right)\:\times\:magnitude\:of\:magnetic\:field\left(B\right)""E=CB"

"E=\left(1.50\times10^{-7}\right)\times\left(3\times10^8\right)=45.0\:N/C"